1) the slope is 2 and the graph passes through ordered pair (1,-5)..
The equation of the line is
(y+5) = 2(x-1)
y+5 = 2x-2
y = 2x-7
why do we add "-5 to y" and why isnt the answer y=2x-5

Answer by Edwin McCravy(20086)

1) the slope is 2 and the graph passes through ordered pair (1,-5)..
The equation of the line is
(y+5) = 2(x-1)
y+5 = 2x-2
y = 2x-7
why do we add "-5 to y" and why isnt the answer y=2x-5


Because we want to solve for y.  When we get to

y + 5 = 2x - 2

we want to leave just the y on the left, so we need
to get rid of the "+ 5" that is just right of the y.

To get rid of that "+ 5"  we put "- 5" under both sides like this:

y + 5 = 2x - 2
  - 5      - 5

Then we draw a line under that:

y + 5 = 2x - 2
  - 5      - 5 

and we then add straight down vertically:

We first bring down the y

y + 5 = 2x - 2
  - 5      - 5
y

Then we combine the "+ 5" and the "- 5" and
get zero, so we don't write anything under that
and then we bring down the = sign:

y + 5 = 2x - 2
  - 5      - 5
y     =

Then we bring down the 2x

y + 5 = 2x - 2
  - 5      - 5
y     = 2x

Finally we combine the "- 2" and the "- 5" and get "- 7"
and write that on the bottom line:

y + 5 = 2x - 2
  - 5      - 5
y     = 2x - 7

Edwin